OpenStudy (anonymous):

ACT math practice question. There are three colors of jelly beans in a bowl. If there are 5 more red ones than blue, and there are 9 more blue ones than green, and the total number of beans is 26, how many green ones are in the bowl? Now I know there should be some equation set up that I can then solve, I just can't think of what the equation would be for some reason. Please help!

OpenStudy (anonymous):

first, create variables to represent the number of each color of jelly beans. For example, let r = number of red jelly beans let b = number of blue jelly beans let g = number of green jelly beans then you can write an equatio for the total, and possibly some other equations. but this gets you started.

OpenStudy (anonymous):

@UsaChan<3 ^

OpenStudy (anonymous):

r=b+5=9+g+5=14+g b=9+g r+b+g=26 find g then others.